### 23 Mar 2026 v1

# thanks: Equal contributionthanks: Equal contribution

Abstract
Inspired by biology’s use of vesicles for targeted transport, many studies have propelled liposomes with active matter, creating synthetic systems that can be viewed as microscale biohybrid robots. Nevertheless, the underlying motility mechanisms from a hydrodynamic perspective are often unresolved, and reliable velocity control remains challenging. Here we present a chlamylipo formed by encapsulating the motile alga Chlamydomonas reinhardtii within a giant liposome. We quantify how the characters of swimming change under controlled perturbations and, from a fluid-mechanical perspective, derive a deformation–velocity expression that incorporates liposome radius, beating frequency, and membrane protrusion. We further show that motility can be reversibly switched by incorporating light-responsive lipids, with the liposome acting as a “clutch” that modulates membrane-coupled propulsion. Thus, liposome encapsulation can function not only as a cargo compartment but also as a tunable motility regulator, enabling speed adjustment and reversible transitions between motile and non-motile states.
preprint: APS/123-QED
Mass transport at cellular and subcellular scales is fundamental to life, and living systems often use vesicles as mobile compartments for material delivery and intercellular signaling. For example, cells transport intracellular vesicles over long distances along microtubules using motor proteins such as dynein and kinesin, internalize and secrete molecular cargo at the cell surface via endocytosis and exocytosis, and release extracellular vesicles that deliver proteins and RNAs to nearby and distant cells; red blood cells, in turn, ferry oxygen through the circulation. To achieve artificial mass transport at the microscale, numerous studies in recent years have focused on the transport of liposomes, synthetic lipid vesicles, driven by active matter. Previous studies have succeeded in transporting liposomes with the external traction force of bacteria [21, 17, 13, 1], sperm [15], myosin [29], and kinesin [6]. These synthetic systems, which combine lipid membranes with living materials, can be regarded as microscale biohybrid robots; such robots integrate living organisms with engineered components to enable novel functionalities and precise control across scales [48].
In contrast to prior studies that relied on external active matter, self-propulsion has also been achieved by encapsulating active matter inside liposomes [23]. In this work, motile Escherichia coli encapsulated in lipid vesicles deform the membrane into tubes that mechanically couple to flagellar motion, turning the tubes into propellers and enabling sustained vesicle swimming. As a related direction beyond liposomes, droplet-based systems have likewise pursued propulsion by encapsulating active matter; theoretical frameworks has been proposed for the encapsulation of a microswimmer [43, 4, 11, 33] or many swimmers [19] inside a droplet.
Figure 1: Typical swimming of chlamylipo. a Schematic illustration of the chlamylipo fabrication method. The water-in-oil emulsion-transfer method was employed to encapsulate the Chlamydomonas into the liposomes. b Long-term swimming behaviour of chlamylipo: the chlamylipo can actively swim even though the lipid membrane encloses the swimmer. The white scale bar represents 20 $\mathrm{\mu m}$. c Time-lapse sequence of a single swimming stroke. During the effective stroke, the Chlamydomonas pushes the liposome surface, causing the membrane protrusion; this non-reciprocal deformation allows the chlamylipo to propel itself. The black scale bar represents 5 $\mathrm{\mu m}$.
Although these biohybrid robots show promise for biomedical applications [9, 26, 51, 10], reliable motion control remains difficult because propulsion is driven by intrinsic biomotor activity with limited tunability. Since motor activity is not readily adjusted by external cues, regulating swimming speed or reversibly switching motility on and off is a nontrivial task. The relation between motor activity and motility from a fluid-mechanical perspective is often lacking, which also makes reliable motion control difficult. In our earlier short communication [45], we proposed a biohybrid robot termed “chlamylipo”, a giant liposome that encapsulates the motile alga Chlamydomonas reinhardtii. The chlamylipo can propel itself even though the liposomal membrane physically isolates the Chlamydomonas from the external world. Because this biohybrid system retains phototaxis, the swimmer’s destination can be steered by the direction of external light. Although we demonstrated that chlamylipo can swim, the underlying swimming mechanism remains unclear. We set two goals in this study. The first goal is to elucidate the swimming mechanism from a fluid mechanics perspective. We aim to derive a simple deformation–velocity relation that predicts motility from observed deformations. The second is to control chlamylipo motility via external stimuli, based on the knowledge of the deformation-velocity relation. In this work, we propose that, beyond cargo transport, liposome encapsulation can function as a motility regulator, allowing speed adjustment and reversible switching between motile and non-motile states.
Result
Persistent swimming of chlamylipo
As illustrated in Fig. 1a, we employed the water-in-oil emulsion-transfer method [35, 36], which is one of the standard methods for fabricating giant unilamellar vesicles (GUVs), to encapsulate Chlamydomonas into liposomes. The chlamylipo was fabricated using the following two steps [45]. First, Chlamydomonas was encapsulated within the water-in-oil emulsion droplets coated with a lipid monolayer. Second, the droplets were centrifuged to pass through the water-oil interface, which was covered with a lipid monolayer, as shown in Fig. 1a, yielding liposomes containing Chlamydomonas. The resulting liposomes have a lipid bilayer membrane; both the interior and exterior phases are aqueous. Note that this emulsion-transfer method has been used to encapsulate other types of active matter, such as the actin filaments [46], actomyosin networks [27, 5], Escherichia coli [31, 23], and bacterial cell division proteins [28, 41]. Compared with other encapsulated systems, motile Chlamydomonas is one of the largest active agents reported to date.
Even though the liposome membrane physically isolates the Chlamydomonas from the surrounding solution, the chlamylipo can achieve persistent swimming [45] as shown in Fig. 1b and Supplemental Movie S1. Same as the free-swimming cells, the chlamylipo swims forward while intermittently reorienting its trajectory, and its swimming velocity was $U=1.43\pm 1.42$ $\mathrm{\mu m/s}$ (79 samples; wild-type strain CC-125 under isotonic outer solution). This velocity is two orders of magnitude slower than the free-swimming cells, which swim with a velocity 100-200 $\mathrm{\mu m/s}$ [30, 25, 14]. Although the swimming strategy of the chlamylipo is not apparent in Supplemental Movie 1 due to its rapid beating, the time-lapsed sequence of a single stroke, shown in Fig. 1c and Supplemental Movie S2, indicates that propulsion arises from shape hysteresis. During the effective stroke (first five images in Fig. 1c), Chlamydomonas pushes the membrane outward, causing the membrane to protrude externally. The position that shows maximum protrusion gradually shifts from the front to the rear during this process. In contrast, during the recovery stroke, the membrane does not demonstrate a significant protrusion, thus contributing to the hysteresis of its deformation. This broken time symmetry in the shape change enables the chlamylipo to achieve the net propulsion [45]. Note that this propulsion strategy, driven by cyclic surface shape deformation, is reminiscent of the envelope model for squirmers [7], and microswimmers in nature such as Volvox [37] and Eutreptiella/Euglena [3].
Figure 2: Deformation-velocity relation of chlamylipo. a Schematic illustrating the variables extracted from image analysis. b Correlation matrix of the measured variables. c Deformation-velocity relation $p^{*}-U^{*}$ of chlamylipo. The experimental conditions were varied by changing the Chlamydomonas strains (wild-type CC-125 and oda1), or by changing the osmotic pressure of the outer solution relative to the inner solution (hypotonic, isotonic, and hypertonic). The inset reports the same $p^{*}-U^{*}$ data, with points color-coded by Chlamydomonas-to-liposome size ratio $\alpha_{\mathrm{CL}}=R_{\mathrm{C}}/R_{\mathrm{L}}$.
Deformation-velocity relation of Chlamylipo
While our earlier short communication [45] described the fabrication protocol and basic swimming behaviour of chlamylipo, a comprehensive physical picture is still lacking. To understand the underlying physics and to establish a simple deformation-velocity equation that is valid over a broad parameter space, we varied the experimental conditions to observe a wide variety of swimming modes. The movements and deformations of chlamylipo are observed under six distinct conditions by modifying the osmotic condition of the outer solution (hypotonic, isotonic, and hypertonic) and using two different Chlamydomonas strains (wildtype CC-125 and oda1). The osmotic condition was varied to alter the reduced volume of the liposome, as discussed in detail later. The flagella-mutant strain oda1 is known for its slower beating of flagella, and its beating frequency and swimming velocity are nearly half of those of WT (CC-125) [30, 14]. A schematic in Fig. 2a represents the obtained parameters via image processing; the swimming velocity $U=2.10\pm 2.55$ $\mathrm{\mu m}/s$, the beating frequency $f=23.6\pm 13.8$ Hz, the liposome radius $R_{\mathrm{L}}=7.73\pm 1.75$ $\mathrm{\mu m}$, length of the major axis of Chlamydomonas $R_{\mathrm{C}}=3.72\pm 0.73$ $\mathrm{\mu m}$, and the protrusion $p=0.90\pm 0.45$ $\mathrm{\mu m}$ are obtained from the entire dataset $N=487$ with all six conditions. The protrusion $p=P-R_{\mathrm{L}}$ is evaluated by subtracting the liposome radius $R_{\mathrm{L}}$ from the farthest distance $P$ from the centroid. The dimensionless protrusion is defined as
$p^{*}=\frac{P-R_{\mathrm{L}}}{R_{\mathrm{L}}}=\frac{p}{R_{\mathrm{L}}},$ (1)
and the value was $p^{*}=0.13\pm 0.07$; the result suggests that the protrusion $p$ is approximately $13\%$ of the liposome radius on average. Now, can parameters describing membrane deformation be used to explain the observed swimming velocity? Figure 2b shows the correlation matrix of the obtained variables. The same figure with the scatter plots is shown in Fig. LABEL:fig:S1. While the velocity $U$ does not correlate with the farthest protrusion $P$, it strongly correlates with the dimensionless protrusion $p^{*}$ (correlation $r=0.66$), weakly with $f$ and $R_{\mathrm{C}}$, and has weakly negative correlation with $R_{\mathrm{L}}$. Notably, as we will discuss later, there is a negative correlation between $p^{*}$ and $R_{\mathrm{L}}$ and a weak negative correlation between $p^{*}$ and $f$.
The correlation matrix shown in Fig. 2b does not readily show a clear picture because the influence of one variable is entangled with that of a different variable. To normalize the difference in the variables, we now construct a dimensionless velocity defined as
$U^{*}$ $=\frac{U}{R_{\mathrm{L}}f}.$ (2)
The swimming velocity can be described by multiplying the beating frequency $f$ and the gained distance per unit cycle $U/f$, and the dimensionless velocity $U^{*}$ corresponds to a gained distance per unit cycle normalized with the liposome radius $R_{\mathrm{L}}$. Figure 2c shows $p^{*}-U^{*}$ plots of six different conditions fall onto a single curve. The dimensionless velocity $U^{*}$ is proportional to the square of the $p^{*}$ under the small deformations ($p^{*}\lesssim 0.1$). The quadratic term becomes the leading-order contribution since a time-reversible first-order deformation cannot produce net propulsion in Stokes flow [40], and this quadratic leading order scaling can be seen in many swimmers in Stokes flow, such as the Taylor’s swimming sheet [47], the envelope model [7, 37], amoeboid swimming [50, 42], and the Najafi–Golestanian three-sphere swimmer [32]. To understand how the scaling is modified under the large deformations, we fitted the entire dataset with a simple function
$U^{*}=\frac{U}{R_{\mathrm{L}}f}=c_{2}p^{*2}+c_{3}p^{*3}$ (3)
and obtained parameters $c_{2}=1.15$ and $c_{3}=-1.91$. The fitted curve, plotted in red in Fig. 2c, indicates that the scaling exponent drops slightly below the quadratic value once the deformation enters the large deformation regime. The negative prefactor in the cubic term $c_{3}<0$ shows that the scaling diminishes at large deformations, consistent with the theoretical prediction [20, 50, 42]. Although theoretical analyses consistently establish a quadratic scaling in the small deformation regime, experimental observations of this scaling remain limited, except for a few examples, including those reported for natural [37] and artificial swimmers [24]. The limited number of observations likely reflects the difficulty of varying the deformation level $p$, which is largely set by body geometry and stroke kinematics and is not easily tuned independently. In this context, our measurements provide one of the first experimental demonstrations of the quadratic scaling with a sample size $n\sim O(10^{2})$, and the present setup offers a practical model platform for probing the deformation-velocity relationship under controlled changes in $p$.
[…]
To summarize the obtained result, the deformation-velocity relation of the chlamylipo can now be described with a simple equation
$U=c_{2}R_{\mathrm{L}}fp^{*2}=\frac{c_{2}fp^{2}}{R_{\mathrm{L}}}$ (4)
and the velocity can be estimated using three deformation variables: the liposome radius $R_{\mathrm{L}}$, the beating frequency $f$, and the protrusion $p$. Note that the higher-order correction term $c_{3}p^{*3}$ is ignored here for simplicity. Up to this point, our discussion has primarily focused on building a deformation-velocity relationship using the “outer” membrane properties ($R_{\mathrm{L}}$, $f$, and $p$), but it is interesting to note here that the “inner” configuration of chlamylipo also affects the swimming velocity. The inset of Fig. 2c shows the same $p^{*}-U^{*}$ relationship, but with points color-coded by the Chlamydomonas-to-liposome size ratio $\alpha_{\mathrm{CL}}=R_{\mathrm{C}}/R_{\mathrm{L}}$, and it shows that both the dimensionless protrusion $p^{*}$ and the dimensionless velocity $U^{*}$ are larger for larger $\alpha_{\mathrm{CL}}$. This relationship can be explained once we recognize that the deformation parameters are dictated by $\alpha_{\mathrm{CL}}$; in other expressions $R_{\mathrm{L}}(\alpha_{\mathrm{CL}})$, $f(\alpha_{\mathrm{CL}})$, $p(\alpha_{\mathrm{CL}})$, and consequently, $U(\alpha_{\mathrm{CL}})$. The protrusion $p^{*}$ has a positive correlation with $\alpha_{\mathrm{CL}}$ (Fig. 2b; correlation $r=0.57$) since the tighter confinements leave less space between the cell and the membrane, allowing Chlamydomonas to displace the membrane further outward. The frequency $f$ has a weak negative correlation r=0.12r=-0.12 with $\alpha_{\mathrm{CL}}$ since tighter confinement distorts the flagellar stroke, elevates hydrodynamic friction, and thus slightly lowers the frequency. The liposome size $R_{\mathrm{L}}$ has a negative correlation r=0.64r=-0.64 with $\alpha_{\mathrm{CL}}$, which is obvious from the definition $R_{\mathrm{L}}=R_{\mathrm{C}}/\alpha_{\mathrm{CL}}$. Taken together, chlamylipo with larger $\alpha_{\mathrm{CL}}$ would protrude further and have smaller hydrodynamic friction due to its small radius $R_{\mathrm{L}}$, thereby swims faster; as a result, the ratio $\alpha_{\mathrm{CL}}$ has a strong positive correlation $r=0.40$ with the velocity $U$. The results indicate that the membrane interior governs its deformation and, in turn, determines swimming speed.
Velocity control via reduced volume
Next, we take a closer look in Fig. 3 how the swimming differs by the conditions, and show that the swimming characteristics of chlamylipo can be tuned from outside. We firstly discuss the effect of the reduced volume by changing the outer solution, and secondly, the effect of Chlamydomonas strains.
The reduced volume $\nu$ ($0<\nu<1$) [12, 18, 44] is a dimensionless parameter representing the ratio of the liposome’s actual volume to that of a sphere with the same surface area, defined as $\nu=6\sqrt{\pi}VA^{-3/2}$ where $V$ and $A$ are the volume and the surface area of liposomes, and this parameter shows the degrees of freedom available for membrane deformation. When the reduced volume approaches unity, the liposome is maximally inflated and constrained to a nearly spherical shape. Conversely, when this parameter is small, the excess surface area relative to the volume allows the liposome to deform into various shapes. Since the membrane deformation and protrusion are key factors for swimming, as revealed in the previous sections, we can now use the reduced volume as a controller to regulate motility. In this work, the reduced volume of chlamylipo is altered by changing the osmotic condition of the outer solution [8]. Figure 3a compares the Chlamydomonas-to-liposome size ratio $\alpha_{\mathrm{CL}}$ under different conditions. By altering the outer medium to a hypertonic/hypotonic solution, the liposome radius $R_{\mathrm{L}}$ decreases/increases since the osmotic pressure difference drives outward/inward water flux across the membrane. Therefore, as shown in the figure, the resultant $\alpha_{\mathrm{CL}}$ becomes larger in a hypertonic outer solution and smaller in a hypotonic one; in other words, $\nu$ is smaller in a hypertonic outer solution and larger in a hypotonic one. Due to this modification, the swimming velocity changes drastically as shown in Fig. 3b; under a hypertonic outer solution, the velocity is 3.32 times that of the isotonic case, whereas it slows to 0.69 of the isotonic case in a hypotonic solution. This velocity difference arises since the dimensionless protrusion $p^{*}$ is drastically different, as shown in Fig. 3c. Under the hypertonic outer solution, Chlamydomonas can easily deform the membrane into various shapes owing to the small reduced volume, and this enhanced deformability leads to larger $p^{*}$ and $U$. On the other hand, under the hypotonic solution Chlamydomonas can only partially deform the nearly spherical membrane, resulting in smaller $p^{*}$ and $U$. Note that the beating frequency $f$ has a decreasing trend with the increase of $\alpha_{\mathrm{CL}}$, as discussed in the previous section. To summarize, the swimming velocity can be modulated by modulating the effective degrees of freedom for the membrane deformation.
Figure 4: Photoswitchable motility by utilizing membrane as a “clutch”. a Schematic figure of the reversible motility photoswitching. The chlamylipo becomes motile with a UV pulse, while it can not swim after the blue light pulse. b Time history of the distance from the initial position $|\mathbf{x}(t)-\mathbf{x}(0)|$ and the swimming velocity $U$ from Supplemental Movie S4. Blue and purple vertical lines show the time of the blue light and UV pulses, respectively. c Time-lapsed snapshots from Supplemental Movie S4. The black scale bar represents 20 $\mathrm{\mu m}$. d Drawing letters “CL” by manipulating the chlamylipo trajectory with lights. The trajectory of the top figure is colored with the elapsed time, while the bottom figure is colored with the swimming velocity. The scatter-marker radius is inversely proportional to the velocity magnitude. The black scale bar represents 50 $\mathrm{\mu m}$.
Secondly, we examine how the velocity differs by Chlamydomonas strains. As mentioned earlier, previous studies reported that both the beating frequency and the resultant swimming velocity of the mutant oda1 are nearly half of those of WT [30, 14]. In agreement with these previous studies, the average beating frequency of WT (based on the data with all three osmotic conditions) was $f_{\mathrm{WT}}=39.6\pm 9.2$ Hz while that of oda1 was $f_{\mathrm{oda1}}=14.8\pm 5.6$ Hz, as shown in Fig. 3d. Note that this beating frequency is slow compared to the bulk swimming, as the frequency is $f_{\mathrm{WT}}=50-70$ Hz and $f_{\mathrm{oda1}}=30-40$ Hz under bulk [30, 14], and it can be hypothesized that the viscous friction of flagella during the effective stroke increased by scratching the liposome membrane. Although it was expected that the swimming velocity $U$ would also be slow for chlamylipo with oda1, no significance with WT-chlamylipo in the velocity was found regardless of the osmotic condition, as shown in Fig. 3b; this result suggests that the swimmer which is slow under the bulk solution, somehow becomes comparable with the faster swimmer once embedded as the chlamylipo. The result appears counterintuitive since the swimming velocity $U$ should be proportional to the beating frequency $f$, as also shown in Eq. (4). This comparable velocity can be explained by the large protrusion $p^{*}$ of oda1, shown in Fig. 3c, compensated for the velocity drop originating from $f_{\mathrm{oda1}}<f_{\mathrm{WT}}$. Two hypotheses can be proposed to account for the enlarged protrusion observed in oda1. The first hypothesis is that the membrane’s relaxation time is so long that the flagellar beating is too rapid to displace the membrane sufficiently. To elucidate the present membrane-flagella interaction, we introduce the following simple spring-damper model. The time evolution of the protrusion $p$ can be modeled with an equation
$\eta\dot{p}+kp=F_{0}\sin(2\pi ft)$ (5)
where $k$ is the spring constant of the protrusion, $\eta$ is the hydrodynamic drag coefficient of the membrane recovery, $F_{0}$ is the amplitude of the flagella force, and $f$ is the beating frequency. By solving the steady-state solution, the amplitude of the protrusion is obtained as
$p_{0}=\frac{F_{0}}{k\sqrt{1+(2\pi f\eta/k)^{2}}}=\frac{F_{0}}{k\sqrt{1+(f\tau)^{2}}},$ (6)
where $\tau=2\pi\eta/k$ is the relaxation time of the protrusion. The equation shows that the protrusion becomes smaller for faster beating frequencies, which is in agreement with the experimental results. Note that a simple toy model for the swimming velocity can be obtained from Eqs. (4) and (6) as follows:
$U=\frac{c_{2}F_{0}^{2}}{R_{\mathrm{L}}k^{2}}\cdot\frac{f}{1+(f\tau)^{2}}$ (7)
and this expression suggests the existence of an optimal beating frequency $f_{\mathrm{opt}}=1/\tau$ that maximizes the velocity $U$. The second hypothesis is based on the differences in the beating pattern. Previous work [16] reports that the oda1 flagellar waveform has lower parabolicity than WT, which may contribute to the large protrusion observed in oda1. A more detailed analysis of flagellum-membrane interactions would be an interesting direction for future work.
Figure 5: Cargo transport and release using chlamylipo. a Schematic figure of cargo transport using chlamylipo. b Snapshots of swimming chlamylipo containing micro-sized beads co-encapsulated with Chlamydomonas as cargo. Images are taken from Supplemental Movie S7. The black scale bar represents 10 $\mathrm{\mu m}$. c Cargo release from chlamylipo upon NIR laser irradiation. Co-encapsulated magnetic beads (i) were expelled as the liposome membrane burst (ii), followed by their diffusion away (iii). Images are taken from Supplemental Movie S8. The black scale bars represent 10 $\mathrm{\mu m}$.
Photoswitchable motility with membrane “clutch”
From the preceding analysis, we find that the protrusion $p$ plays an essential role in determining the swimming velocity $U$, and that the protrusion amplitude can be controlled via the reduced volume $\nu$. Although the osmotic condition can modify the reduced volume, this approach is not a practical method to achieve cyclic regulation of the swimming velocity. To achieve instantaneous control of motility, we introduce the photoswitchable phospholipid AzoPC (azobenzene-containing phosphatidylcholine) into the liposomal membrane [38], enabling light-based control of the reduced volume. AzoPC is a photoswitchable phospholipid that undergoes light-induced photoisomerization: blue-light irradiation stabilizes the rod-like trans state, whereas UV irradiation drives the bent/cone-like cis state, as schematically illustrated in Fig. 4a. As a consequence of this conformational change at the lipid level, the liposome surface area increases upon UV irradiation and decreases upon blue-light irradiation, thereby allowing the reduced volume to be modulated by external light [38, 39, 2]. Indeed, as shown in the bottom images of Fig. 4a and Supplemental Movie S3, the inner Chlamydomonas can protrude the membrane outward after irradiating a UV pulse, while the liposome shape stays nearly spherical after irradiating a blue light pulse. It is worth noting that the Chlamydomonas continues to beat its flagella regardless of the membrane state; the membrane therefore functions as a “clutch” that regulates how much of this internal motion is transmitted to the exterior. This reversible switching of the clutch enables us to control the chlamylipo motility on and off. Note it can be also said that changing the motility level effectively modifies the Péclet number, which quantifies the relative importance of advective swimming to the liposome’s diffusive motion.
Figures 4b–c and Supplemental Movie S4 show the chlamylipo movement under alternating UV and blue light pulses at every 20 s intervals; the irradiating time of each pulse was approximately 500 ms. Note that we used the green light to guide the chlamylipo towards the right side, UV and blue light to switch motility, and red light is used for observation, as Chlamydomonas exhibits no phototactic response to red light. The top panel of Fig. 4b shows the distance from the initial position $|\mathbf{x}(t)-\mathbf{x}(0)|$ where $\mathbf{x}(t)$ is the position at time $t$, while the bottom panel shows the swimming velocity $U$. Although the chlamylipo is in the non-motile state at the initial condition (i), it begins to swim forward quickly after the first UV pulse (ii) with a velocity $\approx 2.0$ $\mathrm{\mu m/s}$, and returns to the non-motile state right after the first blue-light pulse (iii). These motile and non-motile states can be switched reversibly and cyclically, and we have successfully repeated this clutch switching more than seven times, as shown in the figures. These two states enable precise manipulation of the swimmer’s motion; Fig. 4d and Supplemental Movie S5 show that we can trace letters “CL”, the initials of chlamylipo, by steering the position. The top image of Fig 4d shows the trajectory colored with the elapsed time, whereas the bottom image is colored with the swimming velocity, with marker radius inversely proportional to the local velocity magnitude. During tracing, after the swimmer reached each corner of CL letters, the chlamylipo was switched to the non-motile state by blue light; these pauses are indicated by large blue markers. The green light position was then shifted to set the next heading before switching back to the motile state. While steering is possible without the clutch, access to motile and non-motile states allowed precise cornering; pausing at each vertex of the “CL” trajectory enabled accurate reorientation and precise letter formation. Although this tracing is primarily a demonstration, the ability to pause the swimmer in place offers several practical applications and may prove valuable, as discussed in the conclusion. We also note that motility can be switched collectively by broadcasting the switching signal to all chlamylipos, or individually by selectively irradiating only the target chlamylipo, as shown in Fig. LABEL:fig:S2 and Supplemental Movie S6.
Cargo transport with chlamylipo
Finally, we demonstrate that the chlamylipo can transport cargo within the membrane, and also release the inner cargo at an arbitrary position. As illustrated in Fig. 5a, the inner space of the liposome can serve as a compartment for encapsulating desired cargo [1]. The chlamylipo that co-encapsulates cargo can be fabricated by preparing an aqueous solution containing both Chlamydomonas and the target cargo during water-in-oil emulsion formation, and we successfully co-encapsulated micro-sized beads, or even Escherichia coli (data not shown). As shown in Fig. 5b and Supplemental Movie S7, the chlamylipo can propel even when co-encapsulated with 0.5 $\mathrm{\mu m}$ polystyrene beads. The swimming velocity of this chlamylipo is 2.89 $\mathrm{\mu m/s}$, which is comparable to that without cargo. After guiding the chlamylipo to the destination, the inner cargo can be released from the membrane using an NIR (near infrared) laser, as shown in Fig. 5c and Supplemental Movie S8. The snapshot images show a chlamylipo containing 1.0 $\mathrm{\mu m}$ magnetic beads (i) that was irradiated with an NIR laser (ii), leading to the release of the beads, which subsequently diffused away (iii). The magnetic beads were expelled because NIR-induced local heating increased the membrane fluidity, leading to transient pore formation and eventual membrane bursting [34, 49, 22]. Although our approach is constrained by the requirement that the cargos must be harmless to Chlamydomonas, such as micro-sized colloids or benign liquid cargo, we demonstrated that cargos can be co-encapsulated, transported to a target destination, and released on demand by NIR irradiation.
Discussion and Conclusions
In this study, we analysed the swimming of chlamylipo from a fluid-mechanical perspective, identified the factors that govern the swimming velocity, and proposed methods to adjust the velocity based on the understanding. A nondimensional analysis indicated that the speed scales with the beating frequency, the square of the protrusion amplitude, and the inverse of the liposome radius. The ratio of the Chlamydomonas radius to the liposome radius also plays an important role: a larger ratio allows the membrane to protrude farther, leading to faster swimming. Thus, if cargo capacity is not a priority and the goal is simply to obtain a faster chlamylipo, the Chlamydomonas-to-liposome size ratio should be set to a larger value. During the analysis, we also found that the mutant oda1, whose beating frequency and free-swimming velocity are about half those of the wild type, reaches a speed comparable to the wild type when encapsulated; the larger protrusion of oda1 appears to compensate for the lower frequency. The swimming velocity can be modulated by tuning the protrusion level in response to external stimuli, such as osmotic conditions and light switching. When the inner volume of the liposome increases due to the osmotic difference across the membrane, the chlamylipo swims poorly as the membrane has only limited degrees of freedom for deformation. When the inner volume decreases, on the other hand, the chlamylipo can swim faster thanks to the rich deformability of the membrane. By introducing the photoswitchable phospholipid AzoPC into the liposomal membrane, the motility can be reversibly controlled with light, and the membrane works as a “clutch”. The chlamylipo switches to a motile state after UV illumination as the effective membrane area increases, and it returns to a non-motile state after subsequent blue light exposure. Although the light-based control used here was a simple two-state switch, finer continuous tuning should be achievable in future work by more finely controlling the membrane area.
As noted in the introduction, transport at the microscale remains a central challenge, and many studies have explored liposome- or droplet-based cargo delivery using either internally encapsulated active matter or external propulsion by surrounding active matter. A distinctive feature of the present study is that the liposome membrane serves not only as a cargo carrier but also as a clutch that regulates swimming velocity. In microorganism-driven biohybrid robots, previous studies have offered limited means to modulate velocity or to reversibly switch motility off, often relying instead on irreversible approaches such as laser-induced disabling of the microswimmer. In this context, the reversible switching of motility demonstrated here adds important functionality to biohybrid robots: it enables cargo retention without transport, thereby allowing transport and release rates to be tuned; and it improves the precision of manipulation by permitting finer control over the position and motion of the robot. We expect this strategy to be applicable to microrobots of other designs and to guide related work.
The present study introduces the idea of using a membrane as a mechanical element that regulates how internal activity is transmitted to the outside, and this role is not limited to motility switching. More broadly, this concept may be useful for the study of active matter interacting with soft, deformable boundaries, because the mechanical state of the boundary can be externally perturbed in a controlled manner. Such systems may provide a useful platform for investigating the coupling between internal active dynamics and boundary mechanics, while also suggesting engineering routes toward motion control or actuation based on regulated membrane deformation. We expect this concept to be useful for future studies of active matter and its control, as well as for the development of related biohybrid robots.
Acknowledgements.
This work was supported by JSPS KAKENHI (Grants No. 21H05879, No. 21H05891, No. 23K22673, No. 23H04418, No. 23K26040, and No. 23H04430) and the Japan Science and Technology Agency PRESTO (Grant No. JP-MJPR21OA). K. A. acknowledges financial support from the Leave a Nest Grant, incu–be Award. We thank Shinji Deguchi, Daiki Tomioka, Othmane Aouane, Paolo Malgaretti, Taro Toyota, and Kenta Ishimoto for fruitful discussions. We also thank Masafumi Hirono for providing Chlamydomonas and for valuable technical feedback.